### v1

@@ 3.2.5 Mapping clusters to guidelines
After summarization, the keyword representation of the clusters was used to map them to the guidelines. The process involved:
@@
    Comparing cluster embeddings with guideline embeddings using cosine similarity. Clusters were assigned to the guideline entry with the highest similarity, provided there is one with similarity >0.50absent0.50>0.50> 0.50.
@@ 3.3 Experiments and evaluation
3.3.1 Pipeline evaluation
To validate the pipeline, we compared its output with labels from the manually annotated subset. The pipeline was executed on the entire dataset, and results for the manually labelled portion were extracted. This approach ensures fairness, as running the pipeline solely on the labelled subset (1,000 records) would limit its ability to form meaningful clusters due to the restricted data size.
Manual annotations identified 49 distinct labels, with only 13 occurring more than 10 times. For evaluation purposes, these 13 labels were retained for comparison, while the remaining labels were grouped into the other category, which we expect to align with unclustered elements.
The results were evaluated in terms of precision, recall, and f1-score. Metrics were calculated at two levels: at aggregate level to evaluate the overall pipeline performance across all labels, and at per-label level to evaluate performance for individual labels.
@@ 3.3.1 Pipeline evaluation
Additionally, we weighted clinical questions by their corresponding referral counts, incorporating the dataset’s real-world distribution into the analysis.
A second type of evaluation is performed to evaluate the quality of prescription appropriateness labels. In this case, we grouped the 13 labels into four categories: appropriate, inappropriate, likely inappropriate and undetermined. In particular, we marked as likely inappropriate some clusters for which the indication was not completely clear (see Results section for details on this), and as undetermined those that were not covered by the guidelines.
In both cases, we compared our pipeline with other types of text representations/clustering algorithms. We considered clustering results obtained by embeddings from the original Umberto model, by embeddings from an Italian Word2Vec model [44], by TF-IDF representations and by the LDA algorithm. See Appendix B for details.
3.3.2 Analysis of the entire dataset
@@
The pipeline is applied to the entire dataset, with results that provide a complete quantification of referrals for each identified category, together with their appropriateness. Moreover, we present stratified results by year (pre-, during, and post-COVID pandemic), Local Health Authority (LHA), and prescribing physician type (general practitioners vs specialists).
@@ 4 Results
Table 1: Results of clusters mapped to the labels defined by annotations on the manually labelled test set, computed considering clinical questions (CQ) and referrals (REF). P=Precision, R=Recall, F1=F1-Score. The values of all metrics are expressed as percentages.
@@
Table 2: Results of the cluster mapping for each category, on the manually labelled test set, computed for clinical questions (CQ) and referrals (REF). Labels with support <10absent10<10< 10 were grouped with other. P=Precision, R=Recall, F1=F1-Score. P=Precision, R=Recall, F1=F1-Score. The values of all metrics are expressed as percentages.
@@
Other 65.37 67.94 65.68 131 62.56 70.59 66.33 187
4.1 Pipeline evaluation
Table 1 presents the results of the evaluation performed on the manually annotated subset. This evaluation compares the labels assigned to the clustered data with the ones defined by manual annotation. Clusters not mapped to any label or unclustered clinical questions were assigned to the other category.
The pipeline demonstrated superior performance compared to alternative text representations and clustering algorithms, confirming the importance of contextual representations. The metrics were computed for both clinical questions (CQ) and referrals (REF). The referral-level scores were consistently higher due to the influence of clinical questions that occur more frequently, and that seem to be easier to cluster, being less detailed.
Figure 4: Confusion matrix of the clustering assignment to the labels on the manually annotated test set, computed considering the number of clinical questions (A) and the number of referrals (B).
@@ 4.1 Pipeline evaluation
The confusion matrices shown in Figure 4 allow for a more detailed analysis of misclassification errors. Errors in varices surgery are often due to misclassifications as varices, a strictly related cluster. Regarding varices, its precision is mainly affected by misclassifications from venous insufficiency and other. Looking at the specific clinical questions manually labelled as venous insufficiency and classified as varices, we noticed that some presented both diagnoses, making this misclassification reasonable. For the remaining classes, misclassified samples fall mainly into the other group, affecting its precision.
Table 3: Results of clusters mapped to the appropriate/inappropriate/likely inappropriate/undetermined categories on the manually labelled test set, computed for clinical questions (CQ) and referrals (REF). P=Precision, R=Recall, F1=F1-Score. The values of all metrics are expressed as percentages.
P-CQ R-CQ F1-CQ P-REF R-REF F1-REF
Umberto-E3C-Pipeline 89.24 88.27 88.64 93.58 91.52 92.51
Umberto-Base 86.02 80.78 83.13 89.15 84.97 86.55
Word2Vec 80.66 74.29 77.01 85.41 77.54 80.88
TF-IDF 73.51 68.45 67.92 79.44 75.70 74.22
LDA 47.61 48.43 47.83 42.64 46.31 41.31
Table 3 shows the results achieved by grouping labels into the appropriate, inappropriate, likely inappropriate and undetermined categories. The results show higher aggregate performance scores compared to Table 1, indicating that most errors occurred within the same appropriateness category rather than across categories. This suggests that distinguishing between appropriate and inappropriate referrals is simpler, while finer distinctions at the label level are more challenging. A detailed list of all identified clusters, together with their mapping, is reported in Appendix B.
Similar results, including validation on a manually annotated subset, are reported in Appendix D for another type of examination.
4.2 Analysis of the entire dataset
Table 4: Summary of referrals’ clusters, mapped to all labels defined in manual annotations. Green indicates appropriate clusters, red inappropriate clusters, yellow clusters more likely to be inappropriate, and light grey those not mentioned in the guidelines.
@@ 4.2 Analysis of the entire dataset
Venous insufficiency 102,339 20,63
Thrombosis 98,821 19,92
Varices 71,269 14,37
Edema 61,822 12,46
(thrombo)phlebitis 49,537 9,99
Pain 21,131 4,26
Vasculopathy 11,244 2,27
Saphenectomy 8,290 1,67
Varices surgery 7,126 1,44
Diabetes 6,583 1,33
Ulcer 6,272 1,26
Paresthesias 4,856 0,98
Claudicatio 3,570 0,72
Dimer 1,601 0,32
Swelling 1,238 0,25
Respiratory insufficiency 1,060 0,21
Heavy legs 705 0,14
Ischaemia 617 0,12
Stent 612 0,12
Dermatitis 474 0,10
Hypercolesterolaemia 416 0,08
Pregnancy 225 0,05
Dialisis 222 0,04
Surgery 198 0,04
Smoker 193 0,04
Dyspnoea 171 0,03
Tingling 170 0,03
Cold legs 147 0,03
Hypoesthesia 80 0,02
Metabolic syndrome 48 0,01
Bleeding 29 0,01
Other 34,917 7,04
@@
Table 4 extends the analysis on the entire dataset, highlighting the appropriateness labels for each cluster. It can be observed that the distribution of the reasons for the referrals is highly unbalanced, with the first five covering more than 75% of referrals. Varices cluster was assigned to likely inappropriate since this diagnosis is inappropriate unless it is associated with surgery, an element that is often difficult to completely exclude, even for a human reader. Summarizing the table, 34.32% of referrals appear to be appropriate, while 34.07% seems to be not appropriate. The remaining 31.61% referrals can be further divided into 14.37% more likely to be not appropriate, 10.20% that are not mentioned on the guidelines, and 7.04% that cannot be clustered or cannot be mapped to the labels defined by manual annotations. Notably, the other group is limited, while there are clusters for many symptoms and diseases not mentioned in the guidelines.
Figure 6: Prescription appropriateness grouped by LHA (A), type of prescribing physicians (B) and year (C)
Figure 6 reports the level of appropriateness stratified by LHA, type of prescribing physicians and year, providing additional insights into the reasons for inappropriate prescriptions.
@@ 5 Discussion
In this study, we analyzed texts of Italian referrals for prescription appropriateness, with a case study on venous echocolordoppler of lower limbs. By introducing a novel approach that clusters referral texts and maps these clusters to existing guidelines, we demonstrated how it is possible to extract meaningful and actionable insights into the reasons behind such prescriptions.
@@
Our work has several strengths. The analysis encompassed 496,971 referrals, covering an entire Italian region with 10 million inhabitants for three years. This large-scale dataset enabled a comprehensive evaluation of prescription practices. Validation was performed on a manually annotated subset, demonstrating that contextual embedding models significantly outperformed simpler clustering techniques in capturing the semantic nuances of referral texts. Analyzing the results in detail, we observed very high performances for most diseases/symptoms, with only two groups with lower recall. Additionally, results on a secondary dataset (reported in Appendix D) confirmed the generalizability of our methodology.
The analysis of the whole dataset allowed us to have, for the first time, a clear picture of referral motivations for this examination. This is a relevant result for public health authorities, considering the absence of structured recording of this information in the Italian NHS.
Our mapping of referrals to guidelines revealed frequent inappropriate uses, as well as numerous reasons not explicitly covered by current guidelines. While individually rare, these reasons cumulatively represent a non-negligible share of referrals, suggesting the need to expand the scope of guidelines. The ambiguity in referrals grouped under varices, where physicians often omit whether surgery is indicated, together with generic reasons like pain and vasculopathy, also points to the importance of requesting more specific information from referring physicians.
Stratification revealed significant geographical variability, suggesting that targeted interventions are needed in specific Local Health Authorities. General practitioners exhibited higher rates of inappropriate referrals compared to specialists. This might have multiple reasons, including larger patient base, more frequent encounter with less severe cases, but also a lower knowledge and/or agreement with guidelines [45, 46].
Temporal trends showed a decline in inappropriate referrals during the initial COVID-19 wave in 2020, followed by an increase in 2021 as the healthcare system normalized. This aligns with literature suggesting that during the pandemic’s peak, physicians prioritized urgent and necessary examinations to manage resource constraints effectively [47, 48, 49, 50].
Despite its strengths, this study has some limitations. Manual annotation was performed on a subset of data due to the large time required, and annotations were conducted by a single annotator, preventing inter-annotator agreement analysis. While valid for most cases, the focus on a single reason per referral overlooks referrals with multiple reasons, which could be addressed through fuzzy clustering.
Another limitation was the lack of access to patient history, which prevented a precise assessment of the appropriateness of certain referrals, such as those for varices, where the presence of a surgery often could not be verified.
It is also essential to note that this analysis does not aim to state the appropriateness of single referrals, for which a comprehensive analysis of the clinical status and history of the patient would be necessary. Instead, it aims to provide a comprehensive, automatized and accurate picture at the regional level.
In future research, we aim to expand the validation of our pipeline to other types of examinations, with larger manually annotated datasets, ensuring broader applicability across healthcare services.

### v2

@@ 2.2.5 Mapping clusters to guidelines
After summarization, the keyword representation of the clusters is used to map them to the guidelines. The process involves:
@@
    Comparing cluster embeddings with guideline embeddings using cosine similarity. Clusters were assigned to the guideline entry with the highest similarity, provided there is one with similarity >0.50>0.50.
@@ 2.3 Experiments and Evaluation
In this section, we detail the experimental settings and the evaluation procedure for our pipeline, which are summarized in Figure 4.
Figure 4: Summary of the experiments and evaluation procedure
2.3.1 Experimental Setup and Constraints
All experiments were conducted in an offline environment with limited computational resources: 16 GB of RAM, 16 cores Intel Xeon CPU, no GPU. For security and privacy reasons the data could not be move out of this environment. These constraints, commonly present in healthcare settings, influenced several pipeline design decisions, such as splitting the dataset into chunks for clustering, avoiding re-training Tranformer-based models and avoiding the use of generative large language models.
2.3.2 Evaluation Procedure
The pipeline was evaluated by comparing its output with the labels on the manually annotated subsets. Rather than applying the pipeline solely to the labelled records—which would hinder clustering performance due to limited data—we ran the full pipeline on the entire datasets and then extracted results for the annotated subsets. This ensured meaningful cluster formation and fair evaluation.
We performed evaluation at two levels:
  • Disease label evaluation: We assessed the accuracy of the disease labels produced by the pipeline by comparing them to ground-truth labels. To reduce noise, we retained only labels with more than 10 occurrences (1%) in the labelled subset, grouping all others into an Other category. Clusters not mapped to any label or unclustered clinical questions were assigned to the Other category. Metrics were computed at:
    • Aggregate level: to measure overall precision, recall, and F1-score across all labels.
    • Per-label level: to evaluate performance on individual categories.
  • Appropriateness label evaluation: The quality of the pipeline’s assignment of appropriateness labels was evaluated using a four-class scheme: appropriate, inappropriate, likely (in)appropriate, and undetermined. The likely (in)appropriate category included referrals where reasons mentioned in the guidelines were present, but the text lacked complete information to assess the (in)appropriateness. The undetermined category was used for referrals not addressed in the clinical guidelines. Also in this case, we computed metrics and the aggregate and per-label levels.
@@ 2.3.2 Evaluation Procedure
At both levels, we computed the metrics at the single clinical questions level and weighting clinical questions by their corresponding referral counts. The latter allows to incorporating the datasets’ real-world distributions into the analysis.
We also qualitatively show t-SNE 2D projections of the embeddings of the labelled clinical questions. Although this type of visualization has known limitations [45], it remains a common approach in the literature for obtaining a qualitative overview of embedding structures [46].
2.3.3 Hyperparameter Selection and Validation Strategy
Clustering algorithm selection and hyperparameter tuning was carried out exclusively on the ECD dataset. The FEC dataset was reserved for external validation only, ensuring that all performance results reported on FEC represent generalization to unseen data and are free from overfitting to the development set.
2.3.4 Comparative Evaluation
To benchmark our pipeline, we compared it with alternative approaches using different text representations and/or clustering methods. These included:
  • Embeddings from the original UmBERTo model (baseline)
  • Embeddings from an Italian Word2Vec model [47]
  • TF-IDF vector representations
  • Latent Dirichlet Allocation (LDA)
  • Direct matching between embeddings of clinical questions and embeddings of diseases from the guidelines
Additional details are provided in Additional File 1, Section C.
2.3.5 Regional and Temporal Analysis
@@
Finally, we applied the pipeline to the full datasets and reported results at a regional level. This provides a complete quantification of referrals for each identified category, together with their appropriateness. Moreover, we present an enrichment analysis stratifying results by year (pre-, during, and post-COVID pandemic), Local Health Authority (LHA), and prescribing physician type (general practitioners vs. specialists).
@@ 3 Results
Table 1: Summary evaluation on the disease labels
@@
Embedding Match 68.29 58.44 59.21 70.59 61.52 62.09
FEC
Umberto-E3C-Pipeline 90.42 88.78 88.60 93.59 92.70 92.67
Umberto-Base 84.56 75.24 79.63 88.58 79.66 83.88
Word2Vec 81.21 70.55 75.51 82.45 73.86 77.92
TF-IDF 58.42 51.23 54.59 60.41 53.49 56.74
LDA 35.26 41.25 38.02 31.78 38.64 34.88
Embedding Match 75.69 66.62 70.54 77.11 69.85 73.03
00footnotetext: Summary results of mapping to the disease labels defined by manual annotations on the test sets, computed for clinical questions (CQ) and referrals (REF). P=Precision- R=Recall- F1=F1-Score. The values of all metrics are expressed as percentages.
Table 2: Summary evaluation on the appropriateness labels
P-CQ R-CQ F1-CQ P-REF R-REF F1-REF
ECD
Umberto-E3C-Pipeline 89.24 88.27 88.64 93.58 91.52 92.51
Umberto-Base 86.02 80.78 83.13 89.15 84.97 86.55
Word2Vec 80.66 74.29 77.01 85.41 77.54 80.88
TF-IDF 73.51 68.45 67.92 79.44 75.70 74.22
LDA 47.61 48.43 47.83 42.64 46.31 41.31
Embedding Match 74.82 61.90 62.83 77.52 62.29 64.46
FEC
Umberto-E3C-Pipeline 92.05 90.96 90.88 94.66 93.96 94.05
Umberto-Base 87.41 80.55 83.84 89.62 82.16 85.73
Word2Vec 83.24 74.15 78.43 85.77 75.63 80.38
TF-IDF 67.11 57.96 62.20 68.14 60.85 64.29
LDA 38.96 45.62 42.03 35.41 42.93 38.81
Embedding Match 80.07 69.92 74.20 81.15 73.92 77.15
00footnotetext: Summary results of mapping to the appropriate/inappropriate/likely (in)appropriate/undetermined categories on the manually labelled test sets, computed for clinical questions (CQ) and referrals (REF). P=Precision- R=Recall- F1=F1-Score. The values of all metrics are expressed as percentages.
3.1 Pipeline evaluation
Table 1 presents the summary results of the evaluation performed on the manually annotated subsets, at the disease level, comparing different approaches. The pipeline with our fine-tuned Umberto embeddings demonstrates superior performance compared to alternative text representations and clustering algorithms, confirming the importance of contextual representations. The approach with direct matching between embeddings of clinical questions and embeddings of diseases from guidelines exhibits discrete results, but lower than our pipeline, confirming that the clustering procedure act as a useful smoothing step. The metrics are computed for both clinical questions (CQ) and referrals (REF). The referral-level scores are higher for all methods, except LDA, due to the influence of clinical questions that occur more frequently, which seem to be easier to cluster for them, being less detailed.
Table 3: Detailed results of the cluster mapping to the appropriateness labels for our pipeline
P-CQ R-CQ F1-CQ #CQ P-REF R-REF F1-REF # REF
ECD
Appropriate 98.55 93.15 95.77 365 98.97 93.95 96.39 512
Inappropriate 91.29 81.88 86.33 320 96.49 89.58 92.91 921
Likely inappr. 43.28 89.23 58.29 65 84.03 89.89 90.86 628
Undetermined 84.62 79.20 81.82 250 87.22 83.33 85.23 426
FEC
Appropriate 99.36 89.44 94.14 521 99.88 96.96 98.40 2501
Inappropriate 95.35 87.23 91.11 47 95.71 88.16 91.78 76
Likely appr. 98.27 87.18 92.39 195 98.88 90.72 94.62 388
Undetermined 75.24 100.00 85.87 237 84.19 100.00 91.42 591
00footnotetext: Detailed results of the cluster mapping to the appropriateness labels for our pipeline, on the manually labelled test sets, computed for clinical questions (CQ) and referrals (REF). P=Precision, R=Recall, F1=F1-Score. The values of all metrics are expressed as percentages.
Table 2 shows the same results achieved by grouping labels into the appropriate, inappropriate, likely (in)appropriate and undetermined categories. These results show higher aggregate performance scores for ECD, compared to Table 1, indicating that most errors occurred within the same appropriateness category rather than across categories. The results for FEC are also superior, but the improvement is limited. This suggests that distinguishing between appropriate and inappropriate referrals is simpler, while finer distinctions at the label level are more challenging. Table 3 details the metric by appropriateness label, for our pipeline. At the clinical question level, it highlights issues with the likely inappropriate class for ECD, which corresponds directly to the varices class at disease level, and it shows slightly lower results for the undetermined class, but at the referrals level the only relevant difference is with the undetermined class for ECD.
Table 4: Detailed results of the cluster mapping for each disease label for our pipeline on ECD dataset
@@ 3.1 Pipeline evaluation
Other 63.57 67.94 65.68 131 62.56 70.59 66.33 187
00footnotetext: Detailed results of the cluster mapping for each disease for our pipeline, on the manually labelled test set for ECD, computed for clinical questions (CQ) and referrals (REF). Labels with support <10<10 are grouped with Other. P=Precision, R=Recall, F1=F1-Score. The values of all metrics are expressed as percentages.
Table 5: Detailed results of the cluster mapping for each disease label for our pipeline on FEC dataset
Category P-CQ R-CQ F1-CQ # CQ P-REF R-RE F1-REF # REF
Bleeding 100.00 77.91 87.59 163 100.00 97.31 98.64 1748
Polyps 99.35 97.47 98.40 158 99.71 98.55 99.12 344
Colon cancer 96.84 85.98 91.09 107 98.24 89.30 93.56 187
Abdominal pain 88.57 89.42 89.00 104 93.66 89.72 91.65 214
Anaemia 92.75 98.46 95.52 65 95.73 97.39 96.55 115
Constipation 78.38 80.56 79.45 36 91.61 94.24 92.91 139
Diverticulosis 90.00 81.82 85.71 33 86.67 81.25 83.87 48
Altered bowel 84.85 93.33 88.89 30 92.54 95.38 93.94 65
Chronic diarrhoea 100.00 82.14 90.20 28 100.00 93.46 96.62 107
Diverticulitis 100.00 92.86 96.30 14 100.00 89.29 94.34 28
Other 47.14 98.51 63.77 67 59.93 99.42 74.78 173
00footnotetext: Detailed results of the cluster mapping for each disease for our pipeline, on the manually labelled test set for FEC, computed for clinical questions (CQ) and referrals (REF). Labels with support <10<10 are grouped with Other. P=Precision, R=Recall, F1=F1-Score. The values of all metrics are expressed as percentages.
Tables 4 and 5 detail metrics by disease for our pipeline. Most diseases achieve F1-CQ $80\%$, with many exceeding $90\%$. Lower scores are observed for some ECD categories such as varices, varices surgery, vasculopathy, claudicatio and other, while for FEC only other exhibits poor precision. At the referral level (F1-REF), only claudicatio, varices surgery and other achieve scores below $80\%$ for ECD, while for FEC only other is below $80\%$.
Figure 5: Confusion matrix of the clustering assignment to the labels on the manually annotated test sets, for ECD (A, B) and FEC (C,D) computed considering the number of clinical questions (A, C) and the number of referrals (B, D)
The confusion matrices shown in Figure 5 allow for a more detailed analysis of misclassification errors. For ECD, errors in varices surgery are often due to misclassifications as varices, a strictly related cluster. Regarding varices, its precision is mainly affected by misclassifications from venous insufficiency and other. Looking at the specific clinical questions manually labelled as venous insufficiency and classified as varices, we notice that some present both diagnoses, making this misclassification reasonable. For the remaining ECD classes, misclassified samples fall mainly into the other group, affecting its precision. For FEC, the majority of misclassification errors are wrong assignments to the other class. A few are wrong assignments to abdominal pain, which are in multiple cases due to the mention of this symptom along with others that were more relevant. A detailed list of all identified clusters, together with their mapping, is reported in Additional File 1, Section C.
Figure 6 displays t-SNE 2D projections of the embedding vectors of the clinical questions in the two labelled subsets, showing that for most classes there are one or more distinct regions of the 2D space that tend to concentrate the corresponding clinical questions.
Figure 6: 2D t-SNE projections of embeddings of clinical questions from the validation sets of ECD (A) and FEC (B) with their labels, excluding the Other class
3.2 Analysis of the entire dataset
Table 6: Summary of referrals’ clusters for ECD with appropriateness labels
Cluster Label # REF %
@@ 3.2 Analysis of the entire dataset
[…]
00footnotetext: Summary of referrals’ clusters for ECD, mapped to all labels defined in manual annotations. Green indicates appropriate clusters, red inappropriate clusters, yellow clusters for which the information might not be sufficient to define the appropriateness but is more likely inappropriate, and light grey those not mentioned in the guidelines.
Table 7: Summary of referrals’ clusters for FEC with appropriateness labels
@@
[…]
00footnotetext: Summary of referrals’ clusters for FEC, mapped to all labels defined in manual annotations. Green indicates appropriate clusters, red inappropriate clusters, yellow clusters for which the information might not be sufficient to define the appropriateness but is more likely appropriate, and light grey those not mentioned in the guidelines.
Tables 3.2 and 1 extend the analysis to the entire ECD and FEC datasets, highlighting the appropriateness labels for each cluster. We can observe that the distribution of the reasons for the referrals is unbalanced, with 75% of referrals covered by the first five reasons in ECD and by the first seven reasons in FEC. In ECD, varices cluster was assigned to likely inappropriate since this diagnosis is inappropriate unless it is associated with surgery, an element that is often difficult to exclude, even for a human reader. In FEC, colon cancer familiarity / prevention and chronic inflammatory bowel disease were instead marked as likely appropriate since guidelines state that physicians should carefully consider the patient history and condition to establish if and when they should be performed. Specific conditions mentioned in the guidelines — such as requiring ten years from diagnosis or significant colon involvement — were largely absent from referral texts.
Summarizing Table 3.2, for ECD, 34.32% of referrals appear to be appropriate, while 34.07% seems to be not appropriate. The remaining 31.61% referrals can be further divided into 14.37% more likely to be not appropriate, 10.20% that are not mentioned on the guidelines, and 7.04% that cannot be clustered or cannot be mapped to the labels defined by manual annotations. For FEC, 55.98% of referrals appear to be appropriate, while only 5.33% seems to be not appropriate. There are then 15.95% more likely to be appropriate, 15.77% that are not mentioned on the guidelines, and 6.98% that cannot be clustered or cannot be mapped to the labels defined by manual annotations. Notably, in both datasets the other group is limited, while there are clusters for many symptoms and diseases not mentioned in the guidelines, especially for ECD. Although individually rare, these uncovered reasons accounted for a non-negligible proportion of overall referrals, suggesting the need for a broader guideline scope. Additionally, we observed ambiguity in referrals citing varices, where physicians often omitted whether surgical intervention was required. In contrast, FEC referrals showed higher levels of appropriateness, though some inappropriate clusters were identified.
@@
Figure 8: Prescription appropriateness stratified by LHA (A, B), type of prescribing physicians (C, D) and year (E, F) for ECD (A,C,E) and FEC (B,D,F).
Figure 8 reports the level of appropriateness stratified by LHA, type of prescribing physicians and year, providing additional insights into the reasons for inappropriate prescriptions. This stratified analysis reveals significant geographical variability for both examinations and especially for ECD, suggesting that targeted interventions are needed in specific Local Health Authorities. General practitioners exhibited higher rates of inappropriate ECD referrals compared to specialist physicians. This might have multiple reasons, including a larger patient base, more frequent encounters with less severe cases, and a lower knowledge and/or agreement with guidelines [48, 49]. For FEC, the rate of inappropriateness is similar but specialized physicians have more appropriate referrals. An important trend emerges when analyzing the impact of COVID-19 on referral practices. During the early pandemic phase in 2020, inappropriate ECD referrals dropped significantly, likely due to resource constraints prompting physicians to prioritize urgent cases. This pattern is consistent with existing literature, which reports a general reduction in non-urgent medical procedures and a shift toward more conservative referral behaviour during the pandemic [50, 51, 52, 53]. However, this decline was temporary, with inappropriate referrals rebounding in 2021 as healthcare services normalized. A similar, though less pronounced, trend was observed for FEC, possibly due to the already limited number of inappropriate referrals for this examination.
@@ 4 Discussion
In this study, we proposed a methodological pipeline leveraging NLP techniques to assess the appropriateness of diagnostic referrals in the Italian NHS. We illstrated its use through two case studies on venous echocolordoppler of the lower limbs in angiology and flexible endoscope colonoscopy in gastroenterology. By applying an unsupervised approach, we uncovered key trends in referral motivations and guideline adherence, offering the first large-scale assessment of this kind in Italy.
Our work presents several strengths. The pipeline operates in an unsupervised setting, without the need for labelled data, which are costly and time-consuming to produce. This makes it applicable to any examination with established guidelines defined in terms of appropriate and inappropriate reasons for referrals. It is also designed to run on large datasets without requiring high computational resources, lowering the barrier for adoption. In addition to quantifying appropriateness, the pipeline offers a detailed picture of referral motivations, thereby revealing reasons not explicitly covered by existing guidelines.
@@
We demonstrated these capabilities on two population-scale datasets, covering an entire Italian region with 10 million inhabitants for three years. Validation on manually annotated subsets confirmed the robustness of our pipeline and demonstrated that contextual embedding models significantly outperform simpler clustering techniques in capturing the semantic nuances of referral texts. We observed consistently high performances at the appropriate/inappropriate level and for most diseases/symptoms, with only two clusters in ECD with lower recall and one cluster in FEC with lower precision. The excellent results on two different examinations, one of which was not used during pipeline development, confirmed the generalizability of our methodology.
The large-scale analysis enabled, for the first time, a comprehensive view of referral motivations for these examinations. This is a relevant result for public health authorities, given the absence of structured recording of this information in the Italian NHS. The identification of undocumented reasons for referrals highlighted potential gaps in the existing guidelines and the stratified analyses further revealed geographical and temporal patterns and variations across physician types, consistent with findings in the literature and suggestive of targeted interventions to reduce inappropriate referrals. Importantly, the findings directly supported the Lombardy Region in issuing a new resolution extending and reinforcing referral guidelines. This represents one of the first documented cases in Italy where an AI-based analysis of routine healthcare data directly shaped health policy.
This study is a relevant example of a real-world, large-scale application of artificial intelligence in healthcare. Despite a growing body of research, the adoption of AI-based solutions in clinical practice faces significant challenges [54] and remains largely dominated by diagnostic tools in medical imaging [55, 56], particularly in Italy [57]. Our work points to a complementary direction: leveraging large-scale textual data to monitor referral practices. In this way, it illustrates how AI can extend its role beyond diagnosis to support healthcare governance and inform evidence-based policy.
Despite its strengths, this study has some limitations. Manual annotation was limited to a subset of the data due to time constraints, and performed by a single annotator, limiting the possibility of inter-annotator agreement analysis. Our approach also assigns a single primary reason per referral, potentially overlooking multi-reason referrals, which could be better addressed through fuzzy clustering techniques. Moreover, the lack of access to patient histories limited our ability to fully assess clinical appropriateness in certain cases, such as referrals for varices, where prior surgical indication is relevant. It is important to clarify that our pipeline does not aim to assess the appropriateness of individual referrals. Instead, our approach provides a systematic, automated, and large-scale evaluation of prescription practices at the regional level, offering valuable insights for policymakers and healthcare administrators.
In future work, we aim to extend the validation of the pipeline to other examinations and broader settings, including larger annotated datasets and other Italian regions, to further support the adoption of data-driven tools for evaluating referral practices and improving healthcare delivery.